# thanks: Contributed equallythanks: Contributed equally

### v2

@@ Abstract
Collective motion as a flock is a widely observed phenomenon in active matter systems. We report a novel flocking transition mechanism in a system of chemically interacting active colloidal particles sustained purely by chemo-repulsive torques at low to medium densities. The basic requirements to maintain the polar liquid flock are excluded volume repulsions and deterministic long-ranged net repulsive torques. The mechanism we report requires that the time scale individual colloids move a unit length to be dominant with respect to the time they deterministically sense chemicals. This can be equivalently interpreted as pair colloids sliding a minimal unit length before deterministically rotating due to chemical interactions. Switching on the translational repulsive forces renders the flock a crystalline structure. We complement these results with an analysis of a continuum hydrodynamic model, with the transition corresponding to destabilization of the flocking state. Finally, we present simulations for flocking in an agent-based model consisting of local update rules of colloidal pairs that slide and rotate upon collision.
@@ I Introduction
The collective swarming behaviour of interacting agents - also called flocking [1] - has been a topic of sustained interest in the field of non-equilibrium statistical physics. Flocking is also a ubiquitous phenomena in the natural world, widely reported in groups of animals such as birds, fish, and mammals [2]. Paradigmatic models that incorporate short-ranged alignment interactions to study flocking have been extensively studied [3, 4, 5, 1] and continue to be investigated [6, 7, 8, 9, 10]. In the Vicsek model [3], local alignment interactions among the individual agents can lead to transition from a disordered state to one with large-scale, coordinated movement. The flocking transition in Vicsek-like models with short-range interactions has been widely studied, and traveling density bands are observed at the onset of transition [4, 5, 11, 12].
@@
The possibility of attaining an emergent global polar order (flocking) arising purely out of dynamical particle-based models, without any explicit alignment interaction, is a subject of recent interest. Studies on these include those studying the interplay between non-equilibrium phase separation [13] and other collective dynamics [14, 15, 16, 17], those that study the effect of velocity alignment interactions [18, 19, 20], models incorporating attractive interactions [21, 22], history-dependence [23, 24], and those that exclusively study the effect of repulsive forces and torques [25, 26]. Though various possible mechanisms deviating from the Vicsek-like phenomenology have been established, the explicit relevance of the separate force and torque contributions have not been established. For instance, migrating cells that display a global polarity have been known to display chemo-tactic attractive interactions in addition to (among others) either a contact “inhibited” or “attracted” locomotion - akin to a long-ranged torque of either repulsive or attractive nature [27, 28, 22]. Thus, open questions remain on the minimal ingredients and mechanism at play arising from the rotational and translational components separately.
@@
In this paper, we report a novel flocking mechanism for polar liquids where colliding pair particles have to sufficiently slowly rotate in response to chemical gradients before moving a unit distance upon collision. The basic ingredients that are sufficient to produce such a flock are thus (i) short-ranged excluded volume repulsion, and (ii) long-ranged (chemo) repulsive torques. There is no need for any additional (chemo)repulsive translational forces between the particles, nor is there a requisite role for noise in destabilizing the flock. The incorporation of long-ranged translational chemo-repulsive forces between the particles induces instead a crystalline flock (with regular lattice spacing). We demonstrate that the flocking persists even for a chemo-attractive translational force.
@@
The remainder of the paper is organized as follows. In section II, we describe our model to study phoretic interactions of active particles. Our main results are given in section III. In this section, we report our results from particle-based numerical simulations by mapping phase diagrams and describe the properties of the distinct flocking states. Then, we complement the results from particle-based simulation with an analysis of a continuum description in section IV. To make sense of these results further, in section V, we construct an agent-based “sliding model”, where the only interactions are pair colloids that slide and rotate upon contact with each other. Finally, we conclude in section VI by summarizing our main results and comparing it with existing literature, along with suggesting directions for future investigations. In what follows, we describe our model and detail our results.
@@
Figure 1: Polar liquid flocks for $\chi_{t}=0$. (a) Phase diagram in the $(\rho,l_{r}/b)$ plane. Symbols here and throughout the paper denoting a flock (yellow triangle), bands (green circles), and disordered (blue squares). Marked (I)-(III) scenario and transition lines (black and red) are described in the schematic shown in Fig. 2. Snapshots of the respective phases are shown in (b)-(d) respectively (color coding for orientation of each particle is in middle panel). (e) The evolution of the polarization is displayed for the corresponding states (similarly colored). See Movie I [29] for dynamics to obtain CLF starting with a disordered state.
@@
Figure 2: Schematic of the mechanism for formation of fluid flocks. Top-to-bottom: Series of scenarios depicted corresponding to parameters choices labelled in Fig.1a. $l_{-}/b$ is annotated as a dotted red line in (a), while the density dependent $l_{+}/b$ in dashed black line. The middle panel (II) shows the pair collision mechanism that would produce a flock; this occurs is the chemical sensing occurs after sliding a length $l_{sl}^{+}$. Top (I) and bottom (III) panels show scenarios where the chemical response is either too quick or too slow, such that a stable flock cannot be formed
@@ II Model
We consider a set of $N$ chemically interacting polar colloids, where chemical is assumed to be instantaneously deposited on the colloidal surfaces [30, 31, 32, 33]. We model the $i$th active particle as a colloid particle centered at $\mathbf{r}_{i}=(x_{i},y_{i})$, confined to move in two-dimensions, which self-propels with a speed $v_{s}$, along the directions $\mathbf{e}_{i}=(\cos\theta_{i},\,\sin\theta_{i})$. Here $i=1,2,3,\dots,N$ and $\theta_{i}$ is the angle made by the orientation $\mathbf{e}_{i}$ with respect to the positive $x$-axis. The orientation $\mathbf{e}_{i}$ of the $i$th particle, given in terms of the angle $\theta_{i}$, changes due to coupling its dynamics to the dynamics of a phoretic (scalar) field $c$, as we describe below. The position $\mathbf{r}_{i}$ and orientation $\mathbf{e}_{i}$ of the $i$th particle is updated as:.
$\dot{\mathbf{r}}_{i}$ $={\mathbf{V}}_{i},\qquad\qquad{\mathbf{V}}_{i}=v_{s}\mathbf{e}_{i}+\chi_{t}\,\mathbf{J}_{i}+\mu\mathbf{F}_{i}+\bm{\eta}^{t}_{i},$ (1a)
$\dot{\mathbf{e}}_{i}$ $={\mathbf{\Omega}}_{i}\times\mathbf{e}_{i},\qquad\mathbf{\Omega}_{i}=\chi_{r}\left(\mathbf{e}_{i}\times\mathbf{\mathbf{{J}}}_{i}.\right)+\bm{\eta}^{r}_{i}.$ (1b)
For full generality, Eq. (1) is written with both translational and rotational terms, the respective chemical susceptibilities are denoted $\chi_{t}$ and $\chi_{r}$ respectively. Here, $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$ are white noises with zero mean and no temporal correlation. The variance of noises $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$, are respectively, $2D_{t}$ and $2D_{r}$. These noises terms are included for the sake of completion and to demonstrate that our results are robust against weak fluctuations. In this paper, the fluctuations are set to be negligibly small to emphasize that flocks can be formed and destroyed by deterministic causes  [^1] Indeed, the polar order is expected to be destroyed at larger values of fluctuations [3]. In this paper, we choose noise to be less dominant to deterministic effects by choosing Péclet number $\mathrm{Pe}=v_{s}/(bD_{r})$ to be around $10^{9}$.
The inter-particle interaction (of chemical origin) is given in terms of the phoretic current: $\mathbf{\mathbf{\mathbf{{J}}}}_{i}(t)=-\left[\mathbf{\nabla}c(\mathbf{r},t)\right]_{\mathbf{r}=\mathbf{r}_{i}}$, where $c(\mathbf{r},t)$ is the concentration of the chemicals (e.g filled micelles in an oil-emulsion system [35, 23]). The concentration field evolves in time as:
@@
Here, $D_{c}$ is the diffusion coefficient and $c_{0}$ is emission constant of the chemical. The solution of the above can be used to compute $\mathbf{{J}}_{i}$ as:
@@
Here, $r_{ij}=|\mathbf{r}_{ij}|$, with $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$. The term proportional to $\chi_{r}$ in Eq.(1) drives orientational changes (turning away from each other if $\chi_{r}>0$) through interparticle chemical interactions. In this paper, we only consider case of $\chi_{r}>0$ when particles turn away from each other due to chemical interactions [24, 23]. Similarly, the term proportional to $\chi_{t}$ ensures repulsion in the positional dynamics if $\chi_{t}>0$. Here, both $\chi_{r}$ and $\chi_{t}$ are positive (unless otherwise stated) and the system is said to be chemo-repulsive. To preclude overlap of the particles, there is a force on the particles: $\mathbf{F}_{i}=-\boldsymbol{\nabla}_{i}\mathcal{U}$, while, $\mathcal{U}=\sum_{i<j}\mathcal{U}^{e}(\mathbf{r}_{i},\mathbf{r}_{j}).$ With $b$ as the colloidal radius, we choose $\mathcal{U}^{e}$ to be of the form: $\mathcal{U}^{e}=\kappa\left(r_{ij}-2b\right)^{2}$ if $r_{ij}<2b$, while it vanishes otherwise. Here, $\kappa$ is a constant. Simulation details along with the full set of parameters used to generate each figure is given in the Appendix A.
@@ III Results
We first summarize our main results for the case of $\chi_{t}=0$ (no chemorepulsive forces between the particles). In particular, we report here polar fluids that are obtained due to chemo-repulsive torques alone ($\chi_{r}>0$). See Fig.1. These flocks are obtained only if pair colloids move at least a distance ($l_{sl}^{+}$, defined later) before deterministically rotating away due to sensing of chemical gradients. The schematic of the mechanism for polar fluids that we report is presented in Fig. 2. For $\chi_{t}>0$, we obtain crystalline flocks that are obtained via a combination of repulsive torques and forces (Fig. 3(c)); we also find these crystals for (a narrow range of) attractive forces (Fig. 3(b)). For the rest of the paper we denote the former as “Chemorepulsive Liquid Flocks” (CLF), and the later as “Chemorepulsive Crystalline Flocks” (CCF).
Figure 3: (a) Phase diagram in the $(l_{r}/b,l_{t}/b)$ plane for the case of $\chi_{t}>0$. Dashed line refers to lower bound on $\chi_{r}$, solid line for lower bound on $\chi_{r}$ from (5). (b) Similar phase diagram for the case of $\chi_{t}<0$ (extension of (a))). The transition line is extrapolated from (3)(a). (c) Shows a snapshot of the CCF phase (same color scheme as in Fig (1)). See Movie II [29] for dynamics to obtain CCF starting with a disordered state. (d) and (e) compare $g(r,\phi)$ for the case of the CLF and the CCF. Corresponding (f) Pair correlation $g(r)$ plot, in the flocking phase, for CLF and CCF.
@@ III.1 CLFs via long-ranged repulsive torques
We first consider the case of $\chi_{t}=0$, thus repulsions are only short-ranged that avoid overlap of the particles in the system via $\mathbf{F}_{i}$. The phase diagram delimiting polar flocks (or CLF), bands, and disordered states are shown in Figure 1(a). The states are delimited via the mean global polarization of the system at the steady state, $\langle|\mathbf{P}|\rangle_{ss}$, where $\mathbf{P}=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{e}_{i}}{|\mathbf{e}_{i}|}$. On the $(\rho,\chi_{r})$ plane (Fig. 1(a))), we see that low density flocks are sustained, that are eventually destabilized at sufficiently high repulsion and densities. These flocks have evidently a liquid-like spatial structure (Fig. 1(b); see also Fig. 3(d)) [25, 36, 37]. The density field $\rho$ is thus not homogeneous across the system (see Fig. 1 (c); (see also Appendix B)). We note that $\tau={b}/{v_{s}}$ is the propulsion time scale. In addition, we define $\tau_{f}=b^{3}/\chi_{r}$, which sets the time scale for deterministic rotation of the particle due to phoretic interactions.
@@
$l_{r}$ $=\frac{\chi_{r}}{bv_{s}},\qquad\quad l_{t}=\frac{\chi_{t}}{b^{2}v_{s}}.$ (4)
They are, respectively, the deterministic rotation and translational lengths; these quantify the length scales within the system where colloids instantaneously sense long-ranged chemical gradients for the rotational and translational dynamics respectively. From the phase diagram, we see that there are separate length scales that arise as a consequence of the collective many-body dynamics.
To make sense of the above phenomenology, let us construct a sliding time (length) scale, which represents the typical time (length) pair colloids would co-move upon collision before rotating away. The sliding time $\tau_{sl}$, given by $\tau_{sl}=\frac{b^{2}}{v_{s}l_{r}}$. $\tau_{sl}$ is time during which pair colloids slide a distance $l_{sl}=v_{s}\tau_{sl}=\frac{v_{s}b^{3}}{\chi_{r}}$ whilst rotating. Let us denote $l_{-}$ as the smallest length to sustain the CLF (corresponding to the lowest transition line), and $l_{+}$ as the (density dependent) length scale above which the CLF is not sustained. The mechanism for the CLF can then be understood as a competition between $l_{r}$ and $l_{+}$, $l_{-}$; this is illustrated in Fig. 2 (annotated accordingly in the phase diagram of 1(a)). In the moderate density region, we consider a series of transitions from disorder-to-order-to-disordered phases. When $l_{r}<l_{-}$ ( corresponding to $l_{sl}>l_{sl}^{-}$) (bottom), we have a sub-deterministic sensing of the chemicals; here the orientational changes of the colloids are too slow compared to the free movement time, pair colloids slide excessively long, hence local polar order cannot be sustained. Increasing $l_{r}$, in the regime of $l_{-}<l_{r}<l_{+}$ (equivalently $l_{sl}^{+}<l_{sl}<l_{sl}^{-}$) (middle), we have that the chemical response time fast enough such that local polar order is sustained, but still sufficiently slow to break forward-backward symmetry at the pair collision level. Finally for $l_{r}>l_{+}$ (equivalently $l_{sl}<l_{sl}^{+}$), the chemical response is sufficiently quick such that there is no symmetry breaking at the pair collision level. We note that from simulations at $\rho\approx 0.1$ these values we find to be $l_{sl}^{+}\sim 8.3b$ and $l_{sl}^{-}\sim 100b$. We conclude that a pair of colloids have to slide at least a distance $l_{sl}^{+}$, before deterministically turning away, as the underlying mechanism for the CLF.

### v3

@@ Abstract
Collective motion as a flock is a widely observed phenomenon in active matter systems. We report a flocking transition mechanism in a system of chemically interacting active colloidal particles sustained purely by chemo-repulsive torques at low to medium densities. The basic requirements to maintain the polar liquid flock are excluded volume repulsions and deterministic long-ranged net repulsive torques. The mechanism we report requires that the time scale individual colloids move a unit length to be dominant with respect to the time they deterministically sense chemicals. This can be equivalently interpreted as pair colloids sliding a minimal unit length before deterministically rotating due to chemical interactions. Switching on the translational repulsive forces renders the flock a crystalline structure. We complement these results with an analysis of a continuum hydrodynamic model, with the transition corresponding to destabilization of the flocking state.
@@ I Introduction
The collective swarming behaviour of interacting agents - also called flocking [1] - has been a topic of sustained interest in the field of non-equilibrium statistical physics. Flocking is also a ubiquitous phenomena in the natural world, widely reported in groups of animals such as birds, fish, and mammals [2]. Paradigmatic models that incorporate local alignment interactions to study flocking have been extensively studied [3, 4, 5, 1] and continue to be investigated [6, 7, 8, 9]. In the Vicsek model [3], local alignment interactions among the individual agents can lead to transition from a disordered state to one with large-scale, coordinated movement. The flocking transition in Vicsek-like models with short-range interactions has been widely studied, and traveling density bands are observed at the onset of transition [4, 5, 10, 11].
@@
The possibility of attaining an emergent global polar order (flocking) arising purely out of dynamical particle-based models, without any explicit alignment interaction, is a subject of recent interest [12]. Studies on these include those studying the interplay between non-equilibrium phase separation [13] and other collective dynamics [14, 15, 16, 17], those that study the effect of velocity alignment interactions [18, 19, 20], models incorporating attractive interactions [21, 22], history-dependence [23, 24], and those that exclusively study the effect of repulsive forces and torques [25, 26]. Though various possible mechanisms deviating from the Vicsek-like phenomenology have been established, the explicit relevance of the separate force and torque contributions have not been established. For instance, migrating cells that display a global polarity have been known to display chemo-tactic attractive interactions in addition to (among others) either a contact “inhibited” or “attracted” locomotion - akin to a long-ranged torque of either repulsive or attractive nature [27, 28, 22]. Thus, open questions remain on the minimal ingredients and mechanism at play arising from the rotational and translational components separately.
@@
In this paper, we report a minimal mechanism for the formation of polar liquids where colliding pair particles have to sufficiently slowly rotate in response to chemical gradients before moving a unit distance upon collision. The basic ingredients that are sufficient to produce such a flock are thus (i) short-ranged excluded volume repulsion, and (ii) long-ranged (chemo) repulsive torques. There is no need for any additional (chemo)repulsive translational forces between the particles, nor is there a requisite role for noise in destabilizing the flock. Furthermore, the incorporation of long-ranged translational chemo-repulsive forces between the particles induces instead a crystalline flock (with regular lattice spacing). Our work thus adds insights to investigations on attaining a global polar order via dynamical mechanisms (i.e. without any explicit alignment interaction) by providing a minimal mechanism, both for flock formation and destabilization, that, to our best knowledge, has not been reported elsewhere.
@@
The remainder of the paper is organized as follows. In section II, we describe our model to study phoretic (chemical) interactions of active particles. In section III, we present our results from particle-based numerical simulations by mapping phase diagrams and describing the properties of the two distinct flocking states. The two distinct flocking states are termed as: CLF (Chemorepulsive Liquid Flocks), and CCF (Chemorepulsive Crystalline Flocks). Chemorepulsive torques between the particles, which turn them away from each other [15, 23, 26], are essential for the formation of both CLF and CCF states. For the formation of CCF, we require both chemorepulsive force as well as chemorepulsive torques. See below for details. Then, we complement the results from particle-based simulation with a stability analysis of a continuum description in section IV. Finally, we conclude in section V by summarizing our main results and comparing it with existing literature, along with suggesting directions for future investigations.
@@
Figure 1: Polar liquid flocks for $\chi_{t}=0$. (a) Phase diagram in the $(\rho,l_{r}/b)$ plane. Here, $l_{-}/b$ is annotated as a dotted red line, while the density dependent $l_{+}/b$ in dashed black line. Symbols here and throughout the paper denoting a flock (yellow triangle), bands (green circles), and disordered (blue squares). These are distinguished with different values of polar order parameter $\mathbf{P}$: flock ($\mathbf{P}>0.9$), bands ($0.1<\mathbf{P}<0.9$) and disordered ($\mathbf{P}<0.1$). Marked (I)-(III) scenario and transition lines (black and red) are described in the schematic shown in Fig. 2. Snapshots of the respective phases are shown in (b)-(d) respectively (color coding for orientation of each particle is in middle panel). (e) The evolution of the polarization is displayed for the corresponding states (similarly colored). See Movie I [29] for dynamics to obtain CLF starting with a disordered state.
@@
Figure 2: Schematic of the mechanism for formation of fluid flocks. Top-to-bottom: Series of scenarios depicted corresponding to parameters choices labelled in Fig.1a. Top (I) and bottom (III) panels show scenarios where the chemical response is either too slow or too quick, such that a stable flock cannot be formed. The middle panel (II) shows the pair collision mechanism that would produce a flock if the time scale follow: $\tau_{sl}^{+}<\tau_{sl}<\tau_{sl}^{-}$. See text for details.
@@ II Model
We consider a set of $N$ chemically interacting polar colloids. We model the $i$th active particle as a colloid particle centered at $\mathbf{r}_{i}=(x_{i},y_{i})$, confined to move in two-dimensions, which self-propels with a speed $v_{0}$, along the directions $\mathbf{e}_{i}=(\cos\theta_{i},\,\sin\theta_{i})$. Here $i=1,2,3,\dots,N$ and $\theta_{i}$ is the angle made by the orientation $\mathbf{e}_{i}$ with respect to the positive $x$-axis. The orientation $\mathbf{e}_{i}$ of the $i$th particle, given in terms of the angle $\theta_{i}$, changes due to coupling its dynamics to the dynamics of a phoretic (scalar) field $c$, as we describe below. The position $\mathbf{r}_{i}$ and orientation $\mathbf{e}_{i}$ of the $i$th particle is updated as:
$\dot{\mathbf{r}}_{i}={\mathbf{V}}_{i},\qquad\dot{\mathbf{e}}_{i}$ $={\mathbf{\Omega}}_{i}\times\mathbf{e}_{i},$ (1)
Here, the translational velocity $\mathbf{V}_{i}$ and angular velocity $\mathbf{\Omega}_{i}$ of the $i$th particle are given as:
${\mathbf{V}}_{i}$ $=v_{0}\mathbf{e}_{i}+\chi_{t}\,\mathbf{J}_{i}+\mu\mathbf{F}_{i}+\bm{\eta}^{t}_{i},$ (2a)
$\qquad\mathbf{\Omega}_{i}$ $=\chi_{r}\left(\mathbf{e}_{i}\times\mathbf{\mathbf{{J}}}_{i}.\right)+\bm{\eta}^{r}_{i}.$ (2b)
Here $v_{0}$ is the self-propulsion speed of an isolated active particle. The term $\mathbf{J}_{i}$ is chemical flux on the location of the $i$th particle, which gives the inter-particle interaction (of chemical origin). The phoretic flux is given as: $\mathbf{\mathbf{\mathbf{{J}}}}_{i}(t)=-\left[\mathbf{\nabla}c(\mathbf{r},t)\right]_{\mathbf{r}=\mathbf{r}_{i}}$, where $c(\mathbf{r},t)$ is the concentration of the chemicals (e.g filled micelles in an oil-emulsion system [30, 23]). The concentration field evolves in time as:
@@
Here, $D_{c}$ is the diffusion coefficient and $c_{0}$ is emission constant of the chemical. We assume the limit in which chemical field leads to instantaneous interactions between the colloidal surfaces [31, 32, 33, 34]. The solution of the Eq.3 can be used to compute $\mathbf{{J}}_{i}$ as:
@@
Here, $r_{ij}=|\mathbf{r}_{ij}|$, with $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$. The term proportional to $\chi_{r}$ in Eq.(2) drives orientational changes (turning away from each other if $\chi_{r}>0$) through interparticle chemical interactions. In this paper, we only consider case of $\chi_{r}>0$ when particles turn away from each other due to chemical interactions [24, 23]. Similarly, the term proportional to $\chi_{t}$ ensures repulsion in the positional dynamics if $\chi_{t}>0$. Here, both $\chi_{r}$ and $\chi_{t}$ are positive (unless otherwise stated) and the system is said to be chemo-repulsive.
In Eqs.(1)-(2), to preclude overlap of the particles, there is a force on the particles: $\mathbf{F}_{i}=-\boldsymbol{\nabla}_{i}\mathcal{U}$, while, $\mathcal{U}=\sum_{i<j}\mathcal{U}^{e}(\mathbf{r}_{i},\mathbf{r}_{j}).$ With $b$ as the colloidal radius, we choose $\mathcal{U}^{e}$ to be of the form: $\mathcal{U}^{e}=\kappa\left(r_{ij}-2b\right)^{2}$ if $r_{ij}<2b$, while it vanishes otherwise. Here, $\kappa$ is a constant. In Eq.2, $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$ are white noises with zero mean and no temporal correlation. The variance of noises $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$, are respectively, $2D_{t}$ and $2D_{r}$. These noises terms are included for the sake of completion and to demonstrate that our results are robust against weak fluctuations. In this paper, the fluctuations are set to be negligibly small to emphasize that flocks can be formed and destroyed by deterministic causes  [^1] Indeed, the polar order is expected to be destroyed at larger values of fluctuations [3]. In this paper, we choose noise to be less dominant to deterministic effects by choosing Péclet number $\mathrm{Pe}=v_{0}/(bD_{r})$ to be around $10^{9}$. Simulation details along with the full set of parameters used to generate each figure is given in the Appendix A. The dimensions of chemical concentration is: $c]=[\text{L}^{-3}]$, and chemical flux are $\mathbf{J}]=[\text{L}^{-4}]$. Thus, the dimensions of parameters are $\chi_{r}]=[\text{L}^{4}\text{T}^{-1}]$ and $\chi_{t}]=[\text{L}^{5}\text{T}^{-1}]$.
@@ III Results
We first summarize our main results for the case of $\chi_{t}=0$ (no chemorepulsive forces between the particles). In particular, we report here polar fluids that are obtained due to chemo-repulsive torques alone ($\chi_{r}>0$). See Fig.1. These flocks are obtained only if pair colloids move at least a distance ($l_{sl}^{+}$, defined later) before deterministically rotating away due to sensing of chemical gradients. The schematic of the mechanism for polar fluids that we report is presented in Fig. 2. For $\chi_{t}>0$, we obtain crystalline flocks that are obtained via a combination of repulsive torques and forces (Fig. 3(c)). We also find these crystalline flocks for (a narrow range of) attractive forces (Fig. 3(b)). In this paper, we denote the former as “Chemorepulsive Liquid Flocks” (CLF), and the later as “Chemorepulsive Crystalline Flocks” (CCF).
Figure 3: (a) Phase diagram in the $(l_{r}/b,l_{t}/b)$ plane for the case of $\chi_{t}>0$. Dashed line refers to lower bound on $\chi_{r}$, solid line for lower bound on $\chi_{r}$ from Eq. (8). (b) Similar phase diagram for the case of $\chi_{t}<0$ (extension of (a)). The transition line is extrapolated from (3)(a). (c) Shows a snapshot of the CCF phase (same color scheme as in Fig (1)). See Movie II [29] for dynamics to obtain CCF starting with a disordered state. (d) and (e) compare $g(r,\phi)$ for the case of the CLF and the CCF. In (f), a comparison of the flocking phase, for CLF and CCF, is shown for the pair correlation $g(r)$.
@@ III.1 CLFs via long-ranged repulsive torques
We first consider the case of $\chi_{t}=0$, thus repulsions are only short-ranged that avoid overlap of the particles in the system via $\mathbf{F}_{i}$. The phase diagram delimiting polar flocks (or CLF), bands, and disordered states are shown in Figure 1(a). The states are delimited via the mean global polarization of the system at the steady state, $\langle|\mathbf{P}|\rangle_{ss}$, where the polarization $\mathbf{P}$ is defined as the average of sum of the orientation of all the particles:
$\mathbf{P}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i}.$ (5)
On the $(\rho,\chi_{r})$ plane (Fig. 1(a))), we see that low density flocks are sustained, that are eventually destabilized at sufficiently high repulsion and densities. These flocks have evidently a liquid-like spatial structure (Fig. 1(b); see also Fig. 3(d)) [25, 36, 37]. The density field $\rho$ is thus not homogeneous across the system (see Fig. 1(c)). We note that $\tau={b}/{v_{0}}$ is the propulsion time scale. In addition, we define $\tau_{r}=b^{4}/\chi_{r}$, which sets the time scale for deterministic rotation of the particle due to phoretic interactions.
@@
$l_{r}$ $=\frac{b\,\tau}{\tau_{r}}=\frac{\chi_{r}}{b^{2}v_{0}},\qquad\quad l_{t}=\frac{\chi_{t}}{b^{3}v_{0}}.$ (6)
Here, $l_{r}$ is a deterministic length scale constructed by multiplying the radius $b$ of the particle with the ratio of propulsion time scale $\tau$ and time scale $\tau_{r}$ which controls rotations of the particle orientations from chemical interactions between particles. On the other hand, $l_{t}$ is a deterministic translational length, which measures ratio of translation due to chemical interactions between particles and one-body propulsion. From the phase diagram, we see that there are separate length scales that arise as a consequence of the collective many-body dynamics. To this end, let us construct a sliding time scale $\tau_{sl}$ and length scale $l_{sl}$. The time-scale $\tau_{sl}$ is time during which pair colloids slide a distance $l_{sl}$ whilst rotating away. Similarly, $l_{sl}$ represents the typical length the pair colloids would co-move upon collision before rotating away. The sliding time $\tau_{sl}$ and length $l_{sl}$ can be written as:
$\tau_{sl}=\frac{b^{2}}{v_{0}l_{r}},\qquad l_{sl}=v_{0}\tau_{sl}=\frac{v_{0}b^{4}}{\chi_{r}}.$ (7)
We next use these to understand the mechanism of formation of CLF.
Let us denote $l_{-}$ as the smallest value of $l_{r}$ to sustain the CLF (corresponding to the lowest transition line in Fig.1a), and $l_{+}$ as the largest value of $l_{r}$ above which the CLF is not sustained. $l_{+}$ is density dependent, as we describe below. The mechanism for the CLF can then be understood as a competition between $l_{r}$ and $_{+}, ll_{-}$; this is illustrated in Fig. 2 (annotated accordingly in the phase diagram of Fig.1(a)). Equivalently, the mechanism can be understood in terms of sliding time scale $\tau_{sl}$ along with $\tau_{sl}^{-}$ and $\tau_{sl}^{+}$. We now use the schematic of Fig. 2 to explain a series of transitions from disorder-to-order-to-disordered phases. When $l_{r}<l_{-}$ (corresponding to $\tau_{sl}>\tau_{sl}^{-}$), we have a sub-deterministic sensing of the chemicals; here the orientational changes of the colloids are too slow compared to the free movement time, pair colloids slide excessively long, hence local polar order cannot be sustained. See panel (I) of Fig. 2. Increasing $l_{r}$, in the regime of $l_{-}<l_{r}<l_{+}$ (equivalently $\tau_{sl}^{+}<\tau_{sl}<\tau_{sl}^{-}$), we have that the chemical response time fast enough such that local polar order is sustained, but still sufficiently slow such that the forward-backward symmetry at the pair collision level is broken. See panel (II) of Fig. 2. Finally for $l_{r}>l_{+}$ (equivalently $\tau_{sl}<\tau^{+}_{sl}$), the chemical response is sufficiently quick such that there is no symmetry breaking at the pair collision level. See panel (III) of Fig. 2. We note that from simulations at $\rho\approx 0.1$, we find these values to be $l_{sl}^{+}\sim 8.3b$ and $l_{sl}^{-}\sim 100b$. Here $l_{sl}^{+}=v_{0}\tau^{+}_{sl}$ and $l_{sl}^{-}=v_{0}\tau^{-}_{sl}$ (note that in our notation choice $l_{sl}^{-}>l_{sl}^{+}$). We conclude that a pair of colloids have to slide at least a distance $l_{sl}^{+}$, before deterministically turning away, as the underlying mechanism for the CLF.